# Advancing Spatio-Temporal Processing in Spiking Neural Networks through Adaptation

### v1

@@ 4.9 Training and Hyperparameter Search Details for all Tasks
4.12 Details to the Burst Sequence Detection (BSD) task
@@
4.13 Optimization-based feature visualization
@@
The visualization of important input features for the trained network shown in Fig. 6 was obtained using optimization-based feature visualization as defined in erhan2009visualizing. We applied the same sampling method to four different networks, one adLIF and one LIF network trained on BSD and one adLIF and one LIF trained on SHD. For both cases we used the same algorithm, but with different parameters, which are listed in Table 2. For both tasks, we selected network instances that achieved high accuracy. The sampling procedure was performed as follows. First, we initialized a random input example $X^{0}\in\mathbb{R}^{T\times C}$ with $C$ as the input dimension and $T$ the input length (number of time steps), both of the same dimension as the data the network was trained on. Here, each $x^{0}_{t,c}$ was drawn from the uniform distribution $\mathcal{U}(0,1)$. At each iteration $k$, the sample was passed through the network and the loss was calculated akin to the loss function used during training on the specific dataset (SHD or BSD) to obtain the gradient $G_{k}=\nabla_{X}L(X,c)|_{X^{k}}$, with respect to a pre-defined target class $c$. This gradient was normalized to obtain $\Delta X^{k}=\eta\frac{G_{k}}{\zeta_{k}}$ with normalization factor $\zeta_{k}=\max\left\{|G_{k}|_{\text{max}},\epsilon\right\}$ using an $\epsilon=10^{-6}$ for numerical stability, and a step size $\eta$. $|\cdot|_{\text{max}}$ denotes the maximum norm returning the highest absolute value. After the gradient update, we clipped the sample $X^{k+1}$ to the positive range, as the data from both datasets, SHD and BSD, is all-positive. We then applied Gaussian smoothing on the data sample $X^{k+1}$. In the BSD case we applied a $1$D smoothing along the time axis using a Gaussian kernel $g(s)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{s^{2}}{2\sigma^{2}}\right)$ via $X^{k+1}\leftarrow\nu X^{k+1}+\gamma(g*X^{k+1})$. Decay $\nu\in[0,1]$ and smoothing coefficient $\gamma\in[0,1]$ are hyperparameters, * denotes a convolution operation. The $\sigma$ parameter of the Gaussian kernel was linearly decayed from $\sigma_{\text{init}}$ to 00 over the $n_{\text{iter}}$ iterations to decrease the effective regularization applied by the kernel smoothing. For the SHD case we applied a $2$D smoothing to the data sample over both the time dimension and the spatial dimension. For the BSD case, we normalized the sample $X^{k+1}$ with $X^{k+1}\leftarrow X^{k+1}\frac{\mu_{\text{all}}}{\mu_{k+1}}$, where scalar $\mu_{\text{all}}$ is the mean spike rate over all data samples and input neurons in the data, and $\mu_{k+1}$ the mean of sample $X^{k+1}$ over both the time and the input neuron axes. For the SHD setup, we ignored the loss for the first $100$ sequence time steps (compare with grey shaded area in Fig. 6e, whereas for the BSD case we ignored the loss for the first $80\%$ of time steps, which was also done during training of the network on the dataset (see Section 4.9).
@@ 4.14 Hyperparameters
Gradient clip val. $1.5$
Table 3: Global hyperparameters Gradient clip val.: Gradient norm to which the gradient was rescaled if exceeded. FF Init: initialization of feed-forward synaptic weights. Rec Init: Initialization of recurrent synaptic weights.
@@
Table 4: Task- and model-specific hyperparameters lr: learning rate for ADAM optimizer, Ep.: number of training epochs, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), dropout: dropout rate
5 Data Availability
@@ 6 Code Availability
7 Acknowledgements
This work has been supported by the “University SAL Labs” initiative of Silicon Austria Labs (SAL) and its Austrian partner universities for applied fundamental research for electronic based systems, by the European Community’s Horizon 2020 FET-Open Programme, grant number 899265, ADOPD, and by NSF EFRI grant #2318152. We thank Felix Effenberger and Sebastian Otte for valuable feedback on the manuscript.
Declarations
@@ 7 Acknowledgements
Author Contribution
[…]
Appendix A Stability issues of EF-adLIF neurons
@@
In all experiments with SE-adLIF neurons, we used the same parameter range as in Fig. 3a: $a\in[0,120]$, $\tau_{u}\in[5,25]$ ms and $\tau_{w}\in[60,300]$ ms. As shown in the same panel, this parameter range is unstable for EF-adLIF, hence we reduced the range of $a$ to $[0,60]$ for all experiments with EF-adLIF to ensure a decay rate $r<1$ for all parameter configurations. If the larger range of $a\in[0,120]$ is used for EF-adLIF, we can observe instabilities during training, resulting in significant degradation of performance. We report the performance drop on the ECG dataset in Table 5. Within these experiments, we observed exploding gradients during training in many of the runs with the extended range of $a$, which were not present for SE-adLIF with the same parameter range. Partially, these exploding gradients can be alleviated by using gradient clipping, resulting in slightly better performance. This experiment demonstrates the sensitivity of EF-adLIF stability to choices of parameter ranges. Note, that this sensitivity can directly be linked to the relation between neuron parameters and eigenvalues of $\bar{A}_{\mathrm{EF}}$, as shown in Fig. 3.
@@
Appendix B Neural heterogeneity of trained adLIF SNNs
In gerstner2014neuronal, the authors show that adLIF models can, depending on their parametrization, account for many experimentally observed neocortical neuron spiking behaviors. It remained an open question whether these different neuronal dynamics and firing patterns could be beneficial in sequence processing tasks. To that end, we analyzed adLIF networks trained on the SHD task and observed a vast heterogeneity in the evolved neuron parameterizations, see Fig. 8.
@@ Appendix B Neural heterogeneity of trained adLIF SNNs
We categorized the observed neurons based on features derived from their parameters and properties. One first distinction that can be made is whether the neurons exhibit integratory (overdamped, type I) or oscillatory membrane potentials (underdamped, type II), shown in Fig. 8a. We found that most of the neurons ($87\%$) showed underdamped behavior. Most of these underdamped neurons converged to rather slow frequencies of $f<50$ Hz, Fig. 8b. As shown in Fig. 2d and e, these slow-frequency neurons can act as feature detectors for input variations much slower than the spike rates. Another categorization can be made based on the parameter $b$: independent of oscillatory properties, this parameter determines the strength of feed-back from neuron spikes to the adaptation variable $w$. Interestingly, neurons with diverse degrees of such spike-triggered adaptation occured across the entire frequency range (color in Fig. 8b). The distribution of $b$ was rather uniform, with two peaks where we clipped the parameter at the minimum and maximum range, representing weak and strong spike-triggered adaptation respectively. Interestingly, overdamped neurons acted on much longer timescales (see Fig. 8d) than underdamped ones. This figure shows the distribution of the effective time constant $\tau_{\mathrm{eff}}=-\frac{\Delta t}{\ln r}$, which is analogous to the membrane time constant of LIF neurons.
The heterogeneous combinations of the spike-triggered adaptation governed by parameter $b$ and the oscillatory behavior governed by parameter $a$ result in diverse non-linear super- and sub-threshold neuron dynamics (Fig. 8e). Based on the overdamped-underdamped split and the distinction into high and low $b$-values, we selected four neurons from the network for further analysis. These four classes are hence given by overdamped-low-$b$ (OL) neurons, that are equivalent to LIF neurons, overdamped-high-$b$ (OH), underdamped-low-$b$ (UL) and underdamped-high-$b$ (UH) neurons. For example, overdamped neurons with high $b$ exhibit a hyper-polarization, similar to a relative refractory period after emitting a spike (2nd row, 1st column). A second example of an interesting learned behavior is given by underdamped neurons with low spike-triggered adaptation (class UL), which respond with an immediate short burst of action potentials to a positive step in the input current, but with a delayed short burst to a negative step (3rd row, 2nd column). Another intriguing property is the interference of spike-triggered adaptation and oscillations, in which an output spike of an UH-neuron could cancel the membrane potential oscillation (4th row, 3rd column). Such properties don’t exist in LIF neurons (class OL), which only spike in response to sufficient input current, without any complex dynamics. In summary, our analysis revealed that recurrent networks of adLIF neurons exhibit a diverse set of neuronal dynamics when trained for optimal performance on a temporal classification task. This diversity is reminiscent of the diversity of neuronal behaviors found in biological neuronal networks izhikevich2001resonate; gerstner2014neuronal.
Appendix C Additional experiments on spring-mass system
Figure 9: Ablation study for AdLIF networks accurately predict trajectories of dynamical systems. Comparison of LIF and adLIF networks trained with and without recurrent connections. Degradation time denotes the time when the model MSE reaches the MSE of the constant model.
Appendix D Performance and visualized features of networks of neurons with threshold adaptation
We evaluated recurrent spiking networks with a modified variant of the ALIF neuron yinAccurateEfficientTimedomain2021 on the BSD task. For the case of $20$ classes this model achieved a test accuracy of $86.53\pm 1.01\%$ over $10$ runs. For these experiments, we used the same model as in the paper from Yin et al. yinAccurateEfficientTimedomain2021 except for the following modifications: Neuron time constants $\tau_{m}$ and $\tau_{\text{adp}}$ were trained with the reparametrization described in Section Training and Hyperparameter Search Details for all Tasks in Methods and the initialization was the same as for our other experiments (see Table 4). We applied the optimization-based feature visualization from Section Networks of adLIF neurons tune to high-fidelity temporal features to an ALIF network trained on the $20$-class BSD task and show the resulting feature maps in Fig. 10. The hyperparameters for the ALIF model for the BSD task are shown in Table 6.
@@ Appendix D Performance and visualized features of networks of neurons with threshold adaptation
Table 6: Hyperparameters of the ALIF model for the BSD task. lr: learning rate for ADAM optimizer, Ep.: number of training epochs, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), dropout: dropout rate

### v3

@@ Methods
To train spiking neural networks for waveform reconstruction, we considered a spectral and a temporal loss. For the temporal loss, our objective was to maximize the likelihood of the target amplitude $\overset{*}{y}_{\text{w}}[k+1]$ at the $k+1$-th time step, conditioned on the sequence of preceding amplitude values $(x_{\text{w}}[1],\dots,x_{\text{w}}[k])$. This optimization problem can be formalized to
$\arg\max_{\theta}P(\overset{*}{\bm{y}}_{\text{w}};\theta)=\prod^{T-1}_{k=1}p(\overset{*}{y}_{\text{w}}[k+1]|x_{\text{w}}[1],\dots,x_{\text{w}}[k];\theta),$ (88)
where $\theta$ are the model parameters. In line with recent studies (e.g. Wavenet van2016wavenet and sampleRNN mehriSampleRNNUnconditionalEndtoEnd2017), we quantize $\overset{*}{\bm{y}}_{\text{w}}$ into $256$ discrete levels, treating $p$ as a categorical distribution. The output layer of all our simulated networks consisted of $256$ leaky-integrator (LI) neurons representing the logits associated with each of these discrete levels. The target class associated with each time step was defined by the quantization scheme:
$\overset{*}{y}_{\text{q}}[k]=\left\lfloor\frac{f_{A}(\overset{*}{y}_{\text{w}}[k]+1)}{\Delta_{\text{q}}}\right\rceil,$ (89)
where $\left\lfloor x\right\rceil$ denotes the rounding operator, $\Delta_{q}=\frac{2}{2^{8}-1}$ is the discretization interval, and subscript $q$ indicates the quantized, categorical representation. The transformation function $f_{A}(y_{\text{w}}[k])$ represents the A-law non-linearity a_law, defined as:
$f_{A}(x)=\text{sign}(x)\begin{cases}\frac{A|x|}{1+\ln(A)}&\text{if }|x|<\frac{1}{A},\vspace{5pt}\\ \frac{1+\ln(A|x|)}{1+\ln(A)}&\text{if }|x|\geq\frac{1}{A},\end{cases}$ (90)
where $A=86.7$ is a hyperparameter, and $|x|$ denotes the absolute value of $x$. This type of invertible log-space mapping was previously used for audio generation in van2016wavenet and is a common technique used in $8$ bit telephony to improve noise robustness a_law, as it is less sensitive to low amplitude noise while maintaining high precision for significant amplitude magnitudes. The original audio data is encoded with $16$ bit, hence for perfect reconstruction with a categorical distribution this would require $65536$ classes. By using a companding algorithm like the A-law, we limit the effects of reducing our precision to $8$ bit. The temporal loss is finally defined as
$\mathcal{L}_{\text{temp}}(\bm{y}_{\text{q}},\overset{*}{\bm{y}}_{\text{q}})=\text{CrossEntropy}(\bm{y}_{\text{q}},\overset{*}{\bm{y}}_{\text{q}})$ (91)
where $\bm{y}_{\text{q}}[k]=\text{softmax}(\frac{\bm{o}[k]}{\tau})$ of model output $\bm{o}[k]\in\mathbb{R}^{256}$ at time $k$ and $\tau\geq 1$ is a temperature hyperparameter, which we initialized at $\tau=10$ and reduced by a factor of $0.95$ every $2000$ training batches, until a minimum value of $1$.
From this network output $\bm{y}_{\text{q}}[k]$ in the quantized domain we obtained the waveform-amplitude $y_{\text{w}}[k]$ by computing a convex sum over the quantization levels via
$y_{\text{w}}[k]=f^{-1}_{A}(\sum\limits^{256}_{i=1}(i-1)\cdot y^{i}_{\text{q}}[k]\cdot\Delta_{\text{q}}-1).$ (92)
Here, $f^{-1}_{A}$ denotes the inverse of the A-law operation (Eq. (90)) and superscript $i$ denotes the $i$-th entry of the quantized vector. While temporal objectives, such as the one described above, have been used successfully in important studies mehriSampleRNNUnconditionalEndtoEnd2017; van2016wavenet, modern approaches to audio generation and compression generally rely on spectral objectives zeghidourSoundStreamEndtoEndNeural2021; defossez2022high. In this work we found that combining spectral and temporal objectives gives the best results. From the reconstructed waveform $\bm{y}_{\text{w}}$ and the target waveform $\overset{*}{\bm{y}}_{\text{w}}$, we computed a spectral loss based on a multi-resolution Mel-spectral short-term Fourier transform (STFT) loss gritsenko2020spectral defined as
$\mathcal{L}_{\text{w}}=\frac{1}{6}\sum_{k\in\{2^{6},2^{7},\ldots,2^{11}\}}\left(\mathcal{L}^{k}_{lin}+\mathcal{L}^{k}_{log}\right),$ (93)
where
$\mathcal{L}^{k}_{lin}(\bm{y}_{\text{v}},\overset{*}{\bm{y}}_{\text{w}})$ $=\frac{1}{TN}\||\text{STFT}^{k}_{\text{Mel}}(\bm{y}_{\text{v}})|-|\text{STFT}^{k}_{\text{Mel}}(\overset{*}{\bm{y}}_{\text{w}})|\|_{1}$ (94)
$\mathcal{L}^{k}_{log}(\bm{y}_{\text{v}},\overset{*}{\bm{y}}_{\text{w}})$ $=\frac{1}{TN}\|\ln(|\text{STFT}^{k}_{\text{Mel}}(\bm{y}_{\text{v}})|)-\ln(|\text{STFT}^{k}_{\text{Mel}}(\overset{*}{\bm{y}}_{\text{w}})|)\|_{1}.$ (95)
Here, $\text{STFT}^{k}_{\text{Mel}}$ is the unnormalized short-term Fourier transform operator over a $k$-length Hann window using the HTK-variant of the Mel frequency filters. $||\cdot||_{1}$ denotes the $L1$ norm, $T$ corresponds to the number of STFT frames, and $N$ the number of Mel filters ($N=128$). We consider a hop length of $k/4$ and the number of frequencies for the STFT was set to $2048$. The overall loss was defined as
$\mathcal{L}_{\text{total}}\left(\bm{y}_{\text{q}},\overset{*}{\bm{y}}_{\text{w}}\right)=\mathcal{L}_{\text{w}}\left(\bm{y}_{\text{w}},\overset{*}{\bm{y}}_{\text{w}}\right)+\mathcal{L}_{\text{temp}}(\bm{y}_{\text{q}},\overset{*}{\bm{y}}_{\text{q}}).$ (96)
For every sample, during training, we ignored the model output during the first $50$ time steps (i.e. $2$ ms) as burn-in time for the model.
For all simulations (LIF, SE-adLIF, and EF-adLIF), the networks consisted of two fully-connected encoder layers, where only the second layer included recurrent connections. From the second layer, the first $16$ neurons were connected to the decoder network, enforcing the information bottleneck. The decoder network was also a two-layer network, with recurrent connections in both layers, followed by an additional non-recurrent output layer of $256$ leaky integrator (LI) neurons. All remaining hyperparameters are listed in Table 8. For each time step $k$, the network output $\bm{y}[k]\in\mathbb{R}^{256}$ was constructed by passing the output of the LI layer through a softmax.
[…]
$\mathcal{L}_{\text{reg}}(r^{n}_{l})=\begin{cases}g_{l}^{+}\left(r^{n}_{l}-t_{l}^{+}\right)^{2}&\text{if }r^{n}_{l}>t^{+}_{l},\\ g_{l}^{-}\left(r^{n}_{l}-t_{l}^{-}\right)^{2}&\text{if }r^{n}_{l}<t^{-}_{l},\\ 0&\text{otherwise},\end{cases}$ (97)
where $r^{n}_{l}$ is the average number of spikes of neuron $n$ in layer $l$ over time, $t_{l}^{+}$, $t_{l}^{-}$, $g_{l}^{+}$ and $g_{l}^{-}$ are layer-specific hyperparameters according to Table 4. $\mathcal{L}_{\text{reg}}$ was added to the total loss from Eq. (96).
Table 4: Spike regularization parameters for the audio compression task, LIF and AdLIF models *) for the second encoder layer, the spike regularization was only applied to the first $16$ neurons, which constitute the bottleneck and are the only neurons connected to the decoder.
Parameter
Encoder layer 1
Encoder layer 2 *
Decoder layer 1
Decoder layer 2
$t_{l}^{-}$ $0.05$ $0.005$ $0.05$ $0.05$
$t_{l}^{+}$ $0.1$ / $0.5$ (LIF) $0.012$ $0.1$ / $0.5$ (LIF) $0.6$
$g_{l}^{-}$ $10$ $10$ $10$ $10$
$g_{l}^{+}$ $10$ $100$ $10$ $10$
Details for simulations in Fig. 6
[…]
Details for the Burst Sequence Detection (BSD) task
@@
Optimization-based feature visualization
@@
The visualization of important input features for the trained network shown in Fig. 7 was obtained using optimization-based feature visualization as defined in erhan2009visualizing. We applied the same sampling method to four different networks, one adLIF and one LIF network trained on BSD and one adLIF and one LIF trained on SHD. For both cases we used the same algorithm, but with different parameters, which are listed in Table 5. For both tasks, we selected network instances that achieved high accuracy. The sampling procedure was performed as follows. First, we initialized a random input example $X^{0}\in\mathbb{R}^{T\times C}$ with $C$ as the input dimension and $T$ the input length (number of time steps), both of the same dimension as the data the network was trained on. Here, each $x^{0}_{t,c}$ was drawn from the uniform distribution $\mathcal{U}(0,1)$. At each iteration $k$, the sample was passed through the network and the loss was calculated akin to the loss function used during training on the specific dataset (SHD or BSD) to obtain the gradient $G_{k}=\nabla_{X}L(X,c)|_{X^{k}}$, with respect to a pre-defined target class $c$. This gradient was normalized to obtain $\Delta X^{k}=\eta\frac{G_{k}}{\zeta_{k}}$ with normalization factor $\zeta_{k}=\max\left\{|G_{k}|_{\text{max}},\epsilon\right\}$ using an $\epsilon=10^{-6}$ for numerical stability, and a step size $\eta$. $|\cdot|_{\text{max}}$ denotes the maximum norm returning the highest absolute value. After the gradient update, we clipped the sample $X^{k+1}$ to the positive range, as the data from both datasets, SHD and BSD, is all-positive. We then applied Gaussian smoothing on the data sample $X^{k+1}$. In the BSD case we applied a $1$D smoothing along the time axis using a Gaussian kernel $g(s)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{s^{2}}{2\sigma^{2}}\right)$ via $X^{k+1}\leftarrow\nu X^{k+1}+\gamma(g*X^{k+1})$. Decay $\nu\in[0,1]$ and smoothing coefficient $\gamma\in[0,1]$ are hyperparameters, * denotes a convolution operation. The $\sigma$ parameter of the Gaussian kernel was linearly decayed from $\sigma_{\text{init}}$ to 00 over the $n_{\text{iter}}$ iterations to decrease the effective regularization applied by the kernel smoothing. For the SHD case we applied a $2$D smoothing to the data sample over both the time dimension and the spatial dimension. For the BSD case, we normalized the sample $X^{k+1}$ with $X^{k+1}\leftarrow X^{k+1}\frac{\mu_{\text{all}}}{\mu_{k+1}}$, where scalar $\mu_{\text{all}}$ is the mean spike rate over all data samples and input neurons in the data, and $\mu_{k+1}$ the mean of sample $X^{k+1}$ over both the time and the input neuron axes. For the SHD setup, we ignored the loss for the first $100$ sequence time steps (compare with grey shaded area in Fig. 7e, whereas for the BSD case we ignored the loss for the first $80\%$ of time steps, which was also done during training of the network on the dataset (see Section ”Training and Hyperparameter Search Details for all Tasks”).
@@
Table 6: Global hyperparameters FF Init: initialization of feed-forward synaptic weights. Rec Init: Initialization of recurrent synaptic weights.
@@
Table 7: Task- and model-specific hyperparameters lr: learning rate for ADAM optimizer, Ep.: number of training epochs, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), dropout: dropout rate, $q$: coefficient of the reparameterization of $a$ and $b$ (see Eq. (75) and (76)).
@@
Table 8: Hyperparameters for the audio reconstruction task. lr: learning rate for ADAM optimizer, which was reduced by a factor of $0.1$ if the SI-SNR on the validation set did not improve after a training epoch, but not further than $10^{-7}$, Ep.: number of training epochs, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), $q,\hat{a},\hat{b}$: coefficient and clipping ranges of the reparameterization of $a$ and $b$ (see Eq. (75) and (76)) respectively.
model lr
# neurons per layer
q $\hat{a}$ $\hat{b}$
$\alpha$ (SLAYER)
c (SLAYER)
Ep.
[$\tau_{u}^{\text{min}},\tau_{u}^{\text{max}}$]
[$\tau_{w}^{\text{min}},\tau_{w}^{\text{max}}$]
$\tau_{\text{out}}$
batch size
SE-adLIF $5\cdot 10^{-4}$ $300$ $120$ $[0,5]$ [0, 2] $5$ $0.4$ $10$ $[5,25]$ $[30,300]$ $[1,10]$ $128$
EF-adLIF $5\cdot 10^{-4}$ $300$ $20$ $[0,1]$ $[0,2]$ $5$ $0.4$ $10$ $[5,25]$ $[30,300]$ $[1,10]$ $128$
LIF $5\cdot 10^{-4}$ $302$- -- -- $5$ $0.1$ $10$ $[5,100]$- $[1,10]$ $128$
Data Availability
@@
[…]
Acknowledgements
This work has been supported by the “University SAL Labs” initiative of Silicon Austria Labs (SAL) and its Austrian partner universities for applied fundamental research for electronic based systems (M.B., R.F., S.S.), by the European Community’s Horizon 2020 FET-Open Programme, grant number 899265, ADOPD (R.L., M.B., R.F), and by NSF EFRI grant #2318152 (R.L.). This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/COE12] (R.L., M.B.). For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission. We thank Felix Effenberger and Sebastian Otte for valuable feedback on the manuscript.
Author Contributions
Competing Interests
@@ Acknowledgements
Supplementary Notes 1
The SE discretization can be applied to other neuron models with bi-directional feedback between neuron state variables. We demonstrate its benefits on a successful recent neuron model, the balanced harmonic resonate-and-fire (BHRF) neuron higuchi2024balanced. The neuron is described by continuous-time equations of a damped harmonic oscillator as
$\dot{u}$ $=-2bu-\omega^{2}v+I$ (98)
$\dot{v}$ $=u,$ (99)
with neuron states $u$ and $v$, input current $I$, frequency parameter $\omega$, and damping coefficient $d$. In contrast to the adLIF neuron model discussed in the main text, where we can compute the oscillation frequency and decay from the state transition matrix, the BHRF neuron model is directly parameterized by the frequency parameter $\omega$ and damping parameter $d$. The parameter $\omega$ thereby directly defines the natural oscillation frequency (which is the oscillation frequency in the absence of damping). Similarly, the damping parameter $d$ directly defines the rate of decay of the neuron. In the original work by Higuchi et al. higuchi2024balanced, the neuron model is discretised using the Euler Forward method to obtain the following discrete state update equations for discrete time step $k$:
$u[k]$ $=u[k-1]+\delta\left(-bu[k-1]-\omega^{2}v[k-1]+I[k]\right)$ (100)
$v[k]$ $=v[k-1]+\delta u[k-1],$ (101)
where $\delta$ is the discretization time step and $I[k]$ is the time-dependent neuron input. The equations above are simplified by omitting some of the mechanisms from the discrete model proposed by the authors to reduce complexity. By using the same approach as described in the main text (Section Stability analysis of discretized adLIF models), we can calculate the effective oscillation frequency $\omega_{\text{eff}}$ and effective damping coefficient $b_{\text{eff}}$ of the discrete BHRF neuron, and investigate how these two characteristics relate to the neuron parameters $\omega$ and $b$. The effective frequency $\omega_{\text{eff}}$ can thereby be obtained via the same method as described in the main text, and $b_{\text{eff}}$ can be computed from the decay rate $r$, by $b_{\text{eff}}=-\frac{\ln r}{\delta}$. The decay rate $r$ is given by the spectral radius of the state transition matrix of the discretised BHRF neuron (analogous to the method applied to the adLIF model in the main text). The relationship between parameters and effective neuron dynamics is shown in Supplementary Fig. 9a, where we show $b_{\text{eff}}$ (top) and $\omega_{\text{eff}}$ (bottom) with respect to the neuron parameters.
The top panel of Supplementary Fig. 9a reveals that the identity-relationship between $b_{\text{eff}}$ and parameter $b$, which is an inherent property of the continuous-time BHRF neuron from Equations (98) and (99), is disrupted by the Euler-Forward discretization. The same disruption can be observed for $\omega_{\text{eff}}$ and $\omega$ in the bottom panel. As we show below, this disruption can be entirely eliminated, if instead of the Euler-Forward discretization the SE-discretization is applied to the BHRF. We can apply the SE discretization to the BHRF by simply replacing $u^{t-1}$ with $u^{t}$ in Eq. (101), resulting in the update equation
$v[k]=v[k-1]+\delta u[k],$ (102)
which we refer to as SE-BHRF in the following. This minor change in the discretised eqution rescues the identity-mapping between parameter $b$ and effective damping $b_{\text{eff}}$ (Supplementary Fig. 9b, top), as well as the near-identity-mapping between parameter $\omega$ and effective oscillation frequency $\omega_{\text{eff}}$ (Supplementary Fig. 9b, bottom) from the continuous model. Note, that for very low frequencies and sufficiently high damping, a damped harmonic oscillator is overdamped and hence does not oscillate, resulting in the triangle-shaped anomaly region in the top-left corner of the $\omega$-$b$ plot shown in the top row of Fig. 9. Furthermore, it can be observed that the contour lines in the bottom panels of Supplementary Fig. 9 are not exactly vertical. This results from the slight influence of $b$ on $\omega_{\text{eff}}$, a naturally occurring phenomenon in a damped harmonic oscillator. Applying the Euler Forward discretization additionally introduces a region of unstable neuron dynamics, that can be related to our observations of the EF-adLIF model shown in Fig. 3 in the main text. The authors of higuchi2024balanced account for this unstable region by computing $b$ as a relative offset from the divergence boundary (the boundary of the unstable region), but as shown in Supplementary Fig. 9, such region does not exist for the SE-BHRF, rendering additional stability measures such as the relative offset obsolete.
Supplementary Figure 1
Figure 9: Comparion between the Euler-Forward and Symplectic-Euler discretization method for the Balanced Harmonic Resonate-and-fire (BHRF) higuchi2024balanced neuron model. (a) Relationship between the BHRF-parameters $\omega$ and $b$ and the effective damping coefficient $b_{\text{eff}}$ (top panel) and the effective oscillation frequency $\omega_{\mathrm{eff}}$ (bottom panel) of the Euler-Forward-discretised BHRF model (details see text). The solid-colored area denotes the unstable region, where the neuron model is diverging. (b) Same as (a), but for the Symplectic-Euler-discretised BHRF model. Details see text.
Supplementary Notes 2
@@
In the SHD, SSC, and ECG experiments with SE-adLIF neurons, we used the same parameter range as in Fig. 3c: $a\in[0,120]$, $\tau_{u}\in[5,25]$ ms and $\tau_{w}\in[60,300]$ ms. As shown in the same panel, this parameter range is unstable for EF-adLIF, hence we reduced the range of $a$ to $[0,60]$ for all experiments with EF-adLIF to ensure a decay rate $r<1$ for all parameter configurations. If the larger range of $a\in[0,120]$ is used for EF-adLIF, we can observe instabilities during training, resulting in significant degradation of performance. We report the performance drop on the ECG dataset in Supplementary Table 9. Within these experiments, we observed exploding gradients during training in many of the runs with the extended range of $a$, which were not present for SE-adLIF with the same parameter range. Partially, these exploding gradients can be alleviated by using gradient clipping, resulting in slightly better performance. This experiment demonstrates the sensitivity of EF-adLIF stability to choices of parameter ranges. Note, that this sensitivity can directly be linked to the relation between neuron parameters and eigenvalues of $\bar{A}_{\mathrm{EF}}$, as shown in Fig. 3.
Supplementary Table 1
@@
Supplementary Notes 3
In gerstner2014neuronal, the authors show that adLIF models can, depending on their parametrization, account for many experimentally observed neocortical neuron spiking behaviors. It remained an open question whether these different neuronal dynamics and firing patterns could be beneficial in sequence processing tasks. To that end, we analyzed adLIF networks trained on the SHD task and observed a vast heterogeneity in the evolved neuron parameterizations, see Supplementary Fig. 10.
We categorized the observed neurons based on features derived from their parameters and properties. One first distinction that can be made is whether the neurons exhibit integratory (overdamped, type I) or oscillatory membrane potentials (underdamped, type II), shown in Supplementary Fig. 10a. We found that most of the neurons ($87\%$) showed underdamped behavior. Most of these underdamped neurons converged to rather slow frequencies of $f<50$ Hz, Supplementary Fig. 10b. As shown in Fig. 2d and e, these slow-frequency neurons can act as feature detectors for input variations much slower than the spike rates. Another categorization can be made based on the parameter $b$: independent of oscillatory properties, this parameter determines the strength of feed-back from neuron spikes to the adaptation variable $w$. Interestingly, neurons with diverse degrees of such spike-triggered adaptation occured across the entire frequency range (color in Supplementary Fig. 10b). The distribution of $b$ was rather uniform, with two peaks where we clipped the parameter at the minimum and maximum range, representing weak and strong spike-triggered adaptation respectively. Interestingly, overdamped neurons acted on much longer timescales (see Supplementary Fig. 10d) than underdamped ones. This figure shows the distribution of the effective time constant $\tau_{\mathrm{eff}}=-\frac{\Delta t}{\ln r}$, which is analogous to the membrane time constant of LIF neurons.
The heterogeneous combinations of the spike-triggered adaptation governed by parameter $b$ and the oscillatory behavior governed by parameter $a$ result in diverse non-linear super- and sub-threshold neuron dynamics (Supplementary Fig. 10e). Based on the overdamped-underdamped split and the distinction into high and low $b$-values, we selected four neurons from the network for further analysis. These four classes are hence given by overdamped-low-$b$ (OL) neurons, that are equivalent to LIF neurons, overdamped-high-$b$ (OH), underdamped-low-$b$ (UL) and underdamped-high-$b$ (UH) neurons. For example, overdamped neurons with high $b$ exhibit a hyper-polarization, similar to a relative refractory period after emitting a spike (2nd row, 1st column). A second example of an interesting learned behavior is given by underdamped neurons with low spike-triggered adaptation (class UL), which respond with an immediate short burst of action potentials to a positive step in the input current, but with a delayed short burst to a negative step (3rd row, 2nd column). Another intriguing property is the interference of spike-triggered adaptation and oscillations, in which an output spike of an UH-neuron could cancel the membrane potential oscillation (4th row, 3rd column). Such properties don’t exist in LIF neurons (class OL), which only spike in response to sufficient input current, without any complex dynamics. In summary, our analysis revealed that recurrent networks of adLIF neurons exhibit a diverse set of neuronal dynamics when trained for optimal performance on a temporal classification task. This diversity is reminiscent of the diversity of neuronal behaviors found in biological neuronal networks izhikevich2001resonate; gerstner2014neuronal.
Supplementary Figure 2
@@
Supplementary Figure 3
Figure 11: Ablation study for ”Accurate prediction of dynamical systems trajectories” Comparison of LIF and adLIF networks trained with and without recurrent connections. Degradation time denotes the time when the model MSE reaches the MSE of the constant model.
Supplementary Notes 4
We evaluated recurrent spiking networks with a modified variant of the ALIF neuron yinAccurateEfficientTimedomain2021 on the BSD task. For the case of $20$ classes this model achieved a test accuracy of $86.53\pm 1.01\%$ over $10$ runs. For these experiments, we used the same model as in the paper from Yin et al. yinAccurateEfficientTimedomain2021 except for the following modifications: Neuron time constants $\tau_{m}$ and $\tau_{\text{adp}}$ were trained with the reparametrization described in Section ”Training and Hyperparameter Search Details for all Tasks” in ”Methods” and the initialization was the same as for our other experiments (see Table 7). We applied the optimization-based feature visualization from Section ”Networks of adLIF neurons tune to high-fidelity temporal features” to an ALIF network trained on the $20$-class BSD task and show the resulting feature maps in Supplementary Fig. 12. The hyperparameters for the ALIF model for the BSD task are shown in Supplementary Table 10.
Supplementary Figure 4
@@
Table 10: Hyperparameters of the ALIF model for the BSD task. lr: learning rate for ADAM optimizer, Ep.: number of training epochs, S-reg: Coefficient for spike regularization, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), dropout: dropout rate
[…]